OpenStudy (anonymous):

9x - 8y - 3(4y + 6x) Distribute the through the parentheses

OpenStudy (anonymous):

can you do it @compaqangel0205 ??? just distribute

OpenStudy (anonymous):

then add like terms

OpenStudy (anonymous):

umm iam lost no i dont know how to do this i stink at math

OpenStudy (anonymous):

9x-8y-12y-18x=0 add like terms @compaqangel0205

OpenStudy (anonymous):

9x-8y-3(4y+6x) -12y-18x 9x-8y-3-12y-18x 9x-18x=-9x -8y-12y=-20y -9x-20y-3<--------- ANSWER

OpenStudy (anonymous):

just only distribute to remove the parenthesis

OpenStudy (anonymous):

thank you aussy iam going to study that so i can understand the problem

OpenStudy (anonymous):

First you distribute the 3 to the 4y and 6x Then you and or subtract all like terms Then you will get your answer :)

OpenStudy (anonymous):